# A Language-Agnostic Hierarchical LoRA-MoE Architecture for CTC-based Multilingual ASR

### v1

@@ IV-D Ablation Results
In addition, Fig. 3 further compares the LID behavior of mHuBERT-CTC-LIDLoRA and the proposed mHuBERT-CTC-HLoRA. HLoRA produces more diagonal-dominant confusion matrices across all five target languages, leading to an average LID accuracy improvement from 90.1% to 97.9%. This gain is attributed to the single-pass, end-to-end HLoRA design, which jointly optimizes LID and ASR using shared LoRA features, avoiding the frozen and non-adaptive LID stage in two-stage inference and enabling more reliable language-agnostic decoding.
Figure 3: Confusion matrix of LID results on the MLC-SLM dev-set for mHuBERT-CTC-LIDLoRA (left) and the proposed mHuBERT-CTC-HLoRA (right). “OTS” denotes the aggregated predictions of the remaining six source-domain languages in the pretrained mHuBERT-CTC model that are not included in the five target-domain languages.
@@ V Conclusion
This paper presented mHuBERT-CTC-HLoRA, a lightweight and streaming-friendly multilingual ASR framework designed for low-resource, language-agnostic scenarios. By hierarchically decomposing LoRA modules into shared and language-specific components, the proposed method effectively balances cross-lingual representation sharing and language-dependent acoustic modeling. Extensive experiments on the MLC-SLM 2025 benchmark demonstrate that HLoRA consistently outperforms strong two-stage language-agnostic baselines while avoiding external LID supervision. Moreover, the single-pass design significantly improves LID robustness and inference efficiency, making the system more suitable for practical multilingual deployment. Our models and source code are publicly released at https://github.com/zhengyuang7/HLoRA to facilitate reproducibility and future research.

### v2

@@ IV-D Ablation Results
In addition, Fig. 3 further compares the LID behavior of LIDLoRA and the proposed HLoRA. HLoRA produces more diagonal-dominant confusion matrices across all five target languages, leading to an average LID accuracy improvement from 90.1% to 97.9%. This gain is attributed to the single-pass, end-to-end HLoRA design, which jointly optimizes LID and ASR using shared LoRA features, avoiding the frozen and non-adaptive LID stage in two-stage inference and enabling more reliable language-agnostic decoding.
Figure 3: Confusion matrix of LID results on the MLC dev-set for LIDLoRA (left) and the proposed HLoRA (right). “OTS” denotes the aggregated predictions of the remaining six source-domain languages in the pretrained mHuBERT-CTC model that are not included in five target-domain languages.
@@ V Conclusion
This paper presented HLoRA, a lightweight and streaming-friendly multilingual ASR framework designed for low-resource, language-agnostic scenarios. By hierarchically decomposing LoRA modules into shared and language-specific components, the proposed method effectively balances cross-lingual representation sharing and language-dependent acoustic modeling. Extensive experiments on the MLC-SLM 2025 benchmark demonstrate that HLoRA consistently outperforms strong two-stage language-agnostic baselines while avoiding external LID supervision. Moreover, the single-pass design significantly improves LID robustness and inference efficiency, making the system more suitable for practical multilingual deployment. Our models and source code are publicly released at https://github.com/zhengyuang7/HLoRA to facilitate reproducibility and future research.